THE INFRARED SPECTRA OF BF_{3} CATION AND BF_{2}OH CATION TRAPPED IN SOLID NEON

\underline{MARILYN E. JACOX} and WARREN E. THOMPSON, Optical Technology Division, National Institute of Standards and Technology, Gaithersburg, MD 20899-8441.

New, more detailed studies of the photoionization and Penning ionization of BF_{3} trapped in solid neon have confirmed the earlier^{a} infrared spectroscopic identification of BF_{2} and BF_{2} cation and have yielded a revised assignment for the infrared absorptions of BF_{3} cation. The position of the absorption attributed to ν_{3} of that molecule is consistent with the distortion of the ground-state cation from D_{3h} symmetry because of strong vibronic interaction between levels of the B̃ ^{2}E′ state and E′ levels of the X̃ ^{2}A_{2}′ ground state, as predicted by Haller and co-workers.^{b} The facile reaction of BF_{3} with traces of H_{2}O desorbed from the walls of the vacuum system leads to the stabilization of sufficient BF_{2}OH for the identification of two vibrational fundamentals of BF_{2}OH cation.

aM. E. Jacox and W. E. Thompson, J. Chem. Phys. 102, 4747 (1995).

bE. Haller, H. Koppel, L. S. Cederbaum, W. von Niessen, and G. Bieri, J. Chem. Phys. 78, 1359 (1983).